OpenStudy (anonymous):

take the inverse laplace transform of x(s)=(1/(s^2+9))(15.25/(s^2+9.61))

OpenStudy (anonymous):

x(s)=(1/(s^2+9))(15.25/(s^2+9.61)) x(s)=(15.25)/(s^2+9)(s^2+9.61), using partial fraction (15.25)/(s^2+9)(s^2+9.61)=A/(s^2+9) + B(s^2+9.61) 15.25=A(s^2+9.61) + B(s^2+9) 15.25=A s^2 + 9.61A + B s^2 + 9B 15.25=9.61A +9B (1) 0= A + B, A=-B sub this in eq(1) 15.25=-9.61B +9B 15.25=-0.61B B=-25 and A=25 therefore X(s)=25/(s^2+9) -25(s^2+9.61) L^1[X(s)]= (25/3)sin 3t -(25/3.1)sin 3.1 t

OpenStudy (anonymous):

L^-1[X(s)]= (25/3)sin 3t -(25/3.1)sin 3.1 t

OpenStudy (anonymous):

is this related to the differential equation we look at earlier? I know nothing about L aplace transforms but the numbers are similar to the diff. question?

OpenStudy (anonymous):

it seems they are similar...

OpenStudy (anonymous):

X(s)=25/(s^2+9) -25/(s^2+9.61) L^-1[X(s)]= (25/3)sin 3t -(25/3.1)sin 3.1 t but the wolfram answer is different....lol